HIGH RESOLUTION LASER EXCITATION SPECTROSCOPY OF SrCH_{3}, CaBH_{4} AND SrBH_{4}

\underline{M. J. DICK}, Department of Physics, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; P. M. SHERIDAN, J.-G. WANG and P. F. BERNATH, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada.

High resolution laser excitation spectroscopy has been used to record the Ã ^{2}E − X̃ ^{2}A_{1} electronic transition of SrCH_{3} in a laser ablation/molecular jet source. SrCH_{3} was synthesized by the reaction of UV-ablated strontium atoms with a 1% mixture of tetramethyl tin in argon. Transitions arising from the K′=1 ← K′′=0, K′=0 ← K′′=1 and K′=2 ← K′′=1 sub-bands have been observed and assigned. Rotational and fine structure parameters have been determined for the Ã ^{2}E state. An analysis of the spin-orbit and spin-rotation constants indicates that the Ã ^{2}E state does not arise entirely from an atomic orbital of p character. In addition, the Jahn-Teller coupling was found to be negligible in the zero-point vibrational level of the Ã ^{2}E state. The rotational constants were used to estimate a structure for SrCH_{3} in the Ã ^{2}E state. The geometry changes observed between the X̃ ^{2}A_{1} and Ã ^{2}E states of SrCH_{3} are similar to those observed for CaCH_{3}. In addition, the B̃ ^{2}E − X̃ ^{2}A_{1} transitions of CaBH_{4} and SrBH_{4} have been observed. The metal borohydrides were synthesized in a jet source by reaction of UV-ablated metal atoms with a 5% mixture of diborane in argon. A rotational analysis is currently in progress and a comparison of the spectroscopic and structural parameters for both molecules will be presented.